OpenStudy (anonymous):

limx->infinity of (1+a/x)^bx Please look at my work below and tell me were I am messing up! @Calculus1

OpenStudy (anonymous):

Intermediate form type 1^infinity lny=limx->infinity (bx)ln(1+(a/x))

OpenStudy (anonymous):

Intermediate form type infinity * 0 lny = limx->infinity ln(1+(a/x))/(1/bx) or ln(1+(a/x)/(-1bx^-2)

OpenStudy (anonymous):

Assuming that means \[\frac{\ln(1+\frac{a}{x})}{\frac{1}{bx^2}} \] good so far.

OpenStudy (anonymous):

Not x^2 in the bottom, sorry...

OpenStudy (anonymous):

\[e^{ab}\] is my guess. is it right?

OpenStudy (anonymous):

^ yup

OpenStudy (anonymous):

wait, i'm confused, give me a second to think

OpenStudy (anonymous):

Okay so I'm taking derivative of lny= limx->infinity ln(1+a/x)/bx = 1/(1+a/x)/-1(bx)^-2

OpenStudy (anonymous):

\[1/(1+a/x)/-bx^{-2}\]

OpenStudy (anonymous):

And this is where I am stuck. Do I plug in infinity now?

OpenStudy (anonymous):

Not quite. There's a mistake somewhere, it should be \[ \ln(y) = bx \ln(1+\frac{a}{x}) =\frac{\ln(1+\frac{a}{x})}{\frac{1}{bx}}\]

OpenStudy (anonymous):

You need to do l'hospitals now though! I got that^

OpenStudy (anonymous):

????/

OpenStudy (anonymous):

\[ \lim_{x \rightarrow \infty} \frac{\ln(1+\frac{a}{x})}{\frac{1}{bx}} = \lim_{x \rightarrow \infty} \frac{\frac{\frac{-a}{x^2}}{1+\frac{a}{x}}}{\frac{-1}{bx^2}} = \lim_{x \rightarrow \infty} \frac{ab}{1+\frac{a}{x}}=ab\]

OpenStudy (anonymous):

forgot about that chain rule on a/x, damn this is getting messy

OpenStudy (anonymous):

Therefore, \[\lim_{x \rightarrow \infty} (1+\frac{a}{x})^{bx} = e^{ab}\] FYI those of us who can and wish to take the time type in LaTeX on here, which makes our answers somewhat slow but infinitely more readable than regular plain text, so if I or somebody else is typing for awhile, we're just doing this: "lim_{x \rightarrow \infty} (1+\frac{a}{x})^{bx} = e^{ab}" which shows up like this \[lim_{x \rightarrow \infty} (1+\frac{a}{x})^{bx} = e^{ab}\]

OpenStudy (anonymous):

how did you get ab for that ugly numerator?

OpenStudy (anonymous):

you should really do this question with your eyeballs.

OpenStudy (anonymous):

\[\lim_{n\rightarrow \infty}(1+\frac{1}{n})^n=e\] and \[\lim_{n\rightarrow \infty}(1+\frac{x}{n})^n=e^x\]

OpenStudy (anonymous):

What, you mean look it up?

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

I don't understand what you did from this: ((-1/x^2)/(1+a/x))/(-1/bx^2)

OpenStudy (anonymous):

Do you know how to simplify improper fractions? Multiply the top and bottom by -bx^2.

OpenStudy (anonymous):

like, that next step. Which you somehow got ab/1+ax

OpenStudy (anonymous):

Erm, how did you multiply ((-a/x^2)/1+a/x)) by -bx^2??

OpenStudy (anonymous):

Um, \[\frac{\frac{-a}{x^2}*-bx^2}{1+\frac{a}{x}} \]

OpenStudy (anonymous):

Okay, that made it clear^^ Thank you so much. And I'm really bad at simplifying improper fractions.